2. A useful generalization known as the Troutons rule states that, for many non-associated liquids at their normal boiling point, the standard molar entropy of vaporization has a value of about 87 J molK. a) Given that the standard molar enthalpy of vaporization for water is 40.7 kJ mol , calculate the standard molar entropy of vaporization of water. b) Does water obey Troutons rule? Explain. c) The normal boiling point of CC2F2 is -29.79 °C and AH ap 20.2 kJ mol. What is the standard molar entropy of vaporization of CClF2?
Answer #1

According to troutons law Standard Molar entropy = 10.5R = 87J/mol/K

Entropy of vapourization = Enthalpy of vapourization/boiling temperature = 40.7/373 = 0.11 kJ = 110J

Water does not obey Troutons law. This is an exception to the troutons rule. Because of hydrogen bonding in water in the liquid phase entropy is low compared to the gas phase.

Boiling point of CCl2F2 = 243.21K

Standar heat capacity = 20.2 kJ/mole

Standard molar entropy = 20.2/243.21 = 83.055kJ/mole/K